what is the length of the third side of a right triangle if one side is 50 feet and the other side is 50 feet?
Answer by KnightOwlTutor(293)

Use the Pythagorean Formula 'a^2+b^2=c^2 since you have a right Triangle
2500+2500=5000
Take the sq rt. of 50,000=70.71 feet